OpenStudy (anonymous):

A jacket usually sells for $63.00. If the jacket is 30% off, and sales tax is 6%, what is the total price of the jacket, including tax?

OpenStudy (anonymous):

A $66.78 B $46.75 C $47.00

OpenStudy (anonymous):

@SatansFate

OpenStudy (anonymous):

In this case you'll need to take the original price of the jacket, then times it by the factors mentioned. To take 30% off the price, that's equivalent to 70% of the original price. Therefore: \[Price = 63.00 \times 0.7\] Then, we need to add the sales tax on top. In this case we're increasing it by 6%, so it needs to be 106% total. \[Price = (63.00 \times 0.7) \times 1.06 = 46.746\]